### v1

@@ 5 Advection-diffusion-reaction system
where $u:\mathcal{G}_{T}\to\mathbb{R}$ is the concentration, ${\boldsymbol{b}}_{M}$ is the advective velocity field, $\mathbb{A}_{M}$ is the diffusion matrix and $c_{M}>0$ is the reaction constant [52, 39, 42]. As in [39, p.304] we will assume that $(\mathbb{A}_{M})_{ij},\;({\boldsymbol{b}}_{M})_{i},\;\nabla_{M}\cdot{\boldsymbol{b}}_{M},\;c_{M}\in L^{\infty}(M),$ together with ${\boldsymbol{b}}_{M}({\boldsymbol{x}})\in T_{{\boldsymbol{x}}}M,\;\forall{\boldsymbol{x}}\in M$ and the requirement for $\mathbb{A}_{M}$ to be symmetric and to map the tangent space $T_{{\boldsymbol{x}}}M$ onto itself.
Extensive literature exists on finite element discretizations of such equations for what concerns the bulk case, see [46, 45]. For closed surfaces the SFEM discretization of parabolic PDEs in the form of (36) has been comprehensively reviewed in [39]. Convergence estimates for its continuous-in-time formulation using general order polynomials are found in e.g. [42]. For the sake of generality, here we focus on moving surfaces with boundary. The problem reads: Find $u_{h}\in V_{h}(\Gamma_{h})$ such that
(37) $\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+\overline{a}_{h}(u_{h}\,,\,v_{h})+(({\boldsymbol{b}}_{rel}u_{h}-\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}u_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}}\,,\,v_{h})_{\partial\Gamma_{h}}$
@@
(39) $\partial\Gamma_{h,D}\cap\partial\Gamma_{h,N}=\emptyset,\quad\partial\Gamma_{h,D}\cup\partial\Gamma_{h,N}=\partial\Gamma_{h}.$
Nitsche’s boundary penalty method is used in order to impose the inhomogeneous Dirichlet boundary condition $u_{h}(t,{\boldsymbol{x}})|_{\partial\Gamma_{h,D}}=u(t,{\boldsymbol{x}})$ weakly for the diffusive term [87, 45]. For convenience, we define $v^{+}=\max\{v,0\}$ and $v^{-}=\min\{v,0\}$. The resulting bilinear forms are given by
(40) $a_{h}(u_{h}\,,\,v_{h})=$ $~\overline{a}_{h}(u_{h}\,,\,v_{h})-((\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}u_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}}\,,\,v_{h})_{\partial\Gamma_{h,D}}$
@@
(41) $\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+a_{h}(u_{h}\,,\,v_{h})=l_{h}(v_{h}),\;\forall v_{h}\in V_{h}(\Gamma_{h}).$
@@ 5.1 Numerical methods for advection-dominant problems
In biophysical cell models, as introduced in P.1, it is not uncommon to incur advection-dominant ADR equations. In our case, even purely parabolic equations might become advection-dominant due to the introduction of an ALE velocity ${\boldsymbol{v}}^{\mathcal{A}}$. In the present work we employ and extend the continuous interior penalty (CIP) stabilization to cope with the possibly dominant advection regime. The CIP method was proposed and analyzed in [21, 20] and then later extended to the case of stationary surfaces in [23]. Moreover, the CutFEM approach for moving surfaces developed in [53] applies a CIP-type stabilization for advection dominant problems on moving surfaces. We chose CIP for the following reasons:
@@
  • Although only weakly consistent, it has proven to be very successful in time-dependent problems given its commutativity with the time derivative [47];
@@
(42) $\{f\}|_{F}=\frac{1}{2}\left(f_{F}^{+}+f_{F}^{-}\right),\quad{\llbracket f\rrbracket|_{F}}=f_{F}^{+}-f_{F}^{-},$
@@
(43) $s_{h}^{b}\left({u}_{h},{v}_{h}\right):=\gamma_{b}\sum_{F\in\mathcal{F}_{h}^{i}}\phi_{F}h_{F}\left(\llbracket{\boldsymbol{b}_{M_{h}}}\cdot\nabla_{M_{h}}{u}_{h}\rrbracket,\llbracket\boldsymbol{b}_{M_{h}}\cdot\nabla_{M_{h}}{v}_{h}\rrbracket\right)_{F},$
@@
(44) $\phi_{F}=\max(\phi_{T^{+}},\phi_{T^{-}}),\quad\phi_{T^{\pm}}=\min(\beta_{T^{\pm}}^{-1}h_{T^{\pm}},\mu_{T^{\pm}}^{-1}),\quad T^{\pm}\in\mathcal{T}_{h}.$
In the above $\beta_{T}$ is a local velocity scale and $\mu_{T}$ is the reciprocal of a time [47]. If ${\boldsymbol{b}}_{M_{h}}\in L^{\infty}(M_{h})\cap C^{0}(M_{h})$, then 43 can be simplified to
(45) $s_{h}^{b}({u}_{h},{v}_{h}):=\gamma_{b}\sum_{F\in\mathcal{F}_{h}^{i}}\beta_{F}h_{F}^{2}\left(\llbracket\nabla_{M_{h}}{u}_{h}\rrbracket,\llbracket\nabla_{M_{h}}{v}_{h}\rrbracket\right)_{F}.$
@@
(46) $\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+a_{h}(u_{h}\,,\,v_{h})+s_{h}^{b}(u_{h}\,,\,v_{h})=l_{h}(v_{h}),\;\forall v_{h}\in V_{h}(\Gamma_{h}).$
@@ 5.2 Numerical results for the ADR system
(47) $\mathbb{A}(t)=\begin{bmatrix}\cos(t)&-\sin(t)&0\\ \sin(t)&\cos(t)&0\\ 0&0&1\end{bmatrix},\quad B(t)=\begin{bmatrix}0\\ 0\\ 0\end{bmatrix}$.
@@
(48) $\partial_{t}u+\nabla_{\Gamma}\cdot({\boldsymbol{b}}_{\Gamma}u)+c_{\Gamma}u=f,$
@@
${\boldsymbol{b}}_{\Gamma}$ $={\boldsymbol{v}}^{\mathcal{A}}+{\boldsymbol{b}}_{rel}={\boldsymbol{v}}^{\mathcal{A}}+\mathbb{P}_{\Gamma}(2,-{\boldsymbol{x}},0),$
@@
    adrSolver 1: CIP stabilized solver 46.
@@
To further test the effectiveness of the algorithm we now consider an ill-posed problem. Maintaining the half-sphere geometry and the transformation as in 47, we solve the pure transport problem
(50) $\partial_{t}u+\nabla_{\Gamma}({\boldsymbol{b}}_{\Gamma}u)=0,$
@@
(51) ${\boldsymbol{b}}_{\Gamma}={\boldsymbol{v}}^{\mathcal{A}}+{\boldsymbol{b}}_{rel}={\boldsymbol{v}}^{\mathcal{A}}+(z,0,-x)\cdot(1-e^{-10z}),\quad u_{0}=e^{-3(x^{2}+y^{2})}.$
@@
It can be seen that the non-stabilized solver is unstable due to the pure advective regime, leading to oscillations that inevitably effect the whole domain 4(a). adrSolver 1 effectively controls the solution’s gradient thanks to the stabilization, generating a non-oscillating solution, but fails to maintain its positivity, see Figure 5 (right plot). This characteristic is lost along $\partial\Gamma$ where the species encounters the steep boundary layer. adrSolver 2 manages, thanks to the cutoff, to keep the solution positive but fails in maintaining mass-preservation as it can be seen in Figure 5 (left plot), where the relative absolute percentage mass error is plotted. adrSolver 4 successfully manages to deal with the boundary layer, maintain positivity, and avoid oscillations. The final solution is shown in 4(b).
@@ 6 Cahn-Hilliard system as a phase separation model
Phase field modeling plays a central role in biophysical systems to investigate phase transition processes. Common phase field models for stationary surfaces are derived from Allen-Cahn and Cahn-Hilliard dynamics [26, 44, 43, 35], where the motion follows the minimization of an energy functional. We will focus here on the functional
(52) $\mathcal{E}_{CH}(u)=\int_{M(t)}\sigma\Bigl(\underbrace{\frac{\varepsilon|\nabla_{\Gamma}u|^{2}}{2}}_{I}+\underbrace{\frac{1}{\varepsilon}F(u)}_{II}\Bigr),$
@@
(53a) $F_{1}(u)$ $:=\frac{1}{4}((1-u)\log(1-u)+(1+u)\log(1+u))+\frac{1-u^{2}}{2},$
(53b) $F_{2}(u)$ $:=\frac{(1-u^{2})^{2}}{4},$
@@
(54) $F_{1}^{\prime}(u)=\log\Bigl(\frac{1+u}{1-u}\Bigr)-u,\text{ and }F_{2}^{\prime}(u)=u^{3}-u\text{ respectively}.$
$F_{1}$ is a Flory-Huggins-type mixing free energy density whose formulation finds its roots in configurational entropy [59, 48]. It can be seen that $F_{1}$ is only well posed for $u\in(-1,1)$. For this reason, a simpler polynomial potential $F_{2}$ is often used. The advantage of the latter is that it is well defined for values of $u$ outside the interval (1,1)(-1,1).
The extension of the Cahn-Hilliard equation to evolving surfaces has seen recent developments due to its applicability to biological membranes and alloys [105, 41]. The evolving surface Cahn-Hillard equation [24, 25] can be written as
(55) $\begin{cases}\partial^{\bullet}u+(\nabla_{\Gamma}\cdot{\boldsymbol{v}})u=\nabla_{\Gamma}(m\nabla_{\Gamma}w),\\ w=-\sigma\varepsilon\Delta_{\Gamma}u+\frac{\sigma}{\varepsilon}F^{\prime}(u),\end{cases}$
subject to the initial condition $u(0)=u_{0}$ for suitable initial data. $w$ represents the chemical potential and $m\in\mathbb{R}^{+}$ is the mobility. As explained in [77], models for evolving surfaces have similar energy functions to models on stationary surfaces, but important differences arise. Even if the evolution of the systems is driven by the minimization of (52), energy dissipation is no longer satisfied since the velocity is viewed as external force that adds to the system. Analogously, mass preservation depends on the velocity. To keep mass conservation, we must assume $\nabla_{\Gamma}\cdot{\boldsymbol{v}}=0$, which is equivalent to assume inextensibility of the cell membrane, and by consequence area preservation. This assumption also allows to consistently extend this one-phase model to $N$-phase models satisfying the hyperlink condition, as described in [77].
@@ 6.1 Numerical methods for Cahn-Hilliard system
For the discretization of (55), the evolutionary surface finite element method (ESFEM) is one of the most popular methods adopted and the one we consider here [39]. Its advantage lies in simple implementation, low memory consumption, ability to handle complex deformations and potential for efficient parallelization. We refer to [77] for a review of alternative methods. We employ a semi-implicit approach where the fully discrete formulation reads: given $u_{h}^{n}\in V_{h}(\Gamma_{h}^{n})$ find $u^{n+1}_{h},w^{n+1}_{h}\in V_{h}(\Gamma_{h}^{n})$ such that
(56a) $n_{h}\mathopen{}\left(\frac{u^{n+1}_{h}-u_{h}^{n}}{\tau}\,,\,\phi_{h}\mathclose{}\right)+(({\boldsymbol{v}}_{h}-{\boldsymbol{v}}^{\mathcal{A}}_{h})u^{n+1}_{h}\,,\,\nabla_{\Gamma^{n}_{h}}\phi_{h})_{\Gamma^{n}_{h}}+mb_{h}(w^{n+1}_{h}\,,\,\phi_{h})=0,$
(56b) $n_{h}(w^{n+1}_{h}\,,\,\psi_{h})-\sigma\varepsilon b_{h}(u^{n+1}_{h}\,,\,\psi_{h})=\frac{\sigma}{\varepsilon}n_{h}(F^{\prime}(u^{n}_{h})\,,\,\psi_{h}),$
@@
(57) $n_{h}(u_{h}\,,\,\phi_{h})=(u_{h}\,,\,\phi_{h})_{\Gamma_{h}^{n}},\quad b_{h}(u_{h}\,,\,\psi_{h})=(\nabla_{\Gamma_{h}^{n}}u_{h}\,,\,\nabla_{\Gamma_{h}^{n}}\psi_{h})_{\Gamma_{h}^{n}}.$
Similar discretizations can be found in [43, 7, 85], where a Cahn-Hilliard solver is coupled with a domain evolution which is unknown. In those works, the potential $F_{2}$ was used. In our case, we can take advantage of the scheme presented in Section 3 to employ both $F_{1}$ and $F_{2}$, and apply mass conservation if needed. The scheme proceeds as follows.
In the predictor step we solve the unconstrained problem 56. Taking the non-linear term explicitly ensures $u_{h}^{n}$ lays inside the desired bound.
@@
Remark 6.1.
Another strength of the postprocessing technique of [29] evident here is that it can act directly on the quantity of interest per se without involving $w^{n+1}_{h}$, which is unbounded. We remind the reader that extensions for higher order BDFs methods are presented in the original articles.

### v2

@@ 5 Advection-diffusion-reaction system
where $u:\mathcal{G}_{T}\to\mathbb{R}$ is the concentration, ${\boldsymbol{b}}_{M}$ is the advective velocity field, $\mathbb{A}_{M}$ is the diffusion matrix and $c_{M}>0$ is the reaction constant Gastaldi [2001], Dziuk and Elliott [2013], Elliott and Ranner [2021]. As in [Dziuk and Elliott, 2013, p.304] it is assumed that $(\mathbb{A}_{M})_{ij},\;({\boldsymbol{b}}_{M})_{i},\;\nabla_{M}\cdot{\boldsymbol{b}}_{M},\;c_{M}\in L^{\infty}(M),$ together with ${\boldsymbol{b}}_{M}({\boldsymbol{x}})\in T_{{\boldsymbol{x}}}M,\;\forall{\boldsymbol{x}}\in M$ and the requirement for $\mathbb{A}_{M}$ to be symmetric and to map the tangent space $T_{{\boldsymbol{x}}}M$ onto itself.
Extensive literature exists on finite element discretizations of such equations for what concerns the bulk case, see Ern and Guermond [2021b, a]. For closed surfaces the SFEM discretization of parabolic PDEs in the form of (37) has been comprehensively reviewed in Dziuk and Elliott [2013]. Convergence estimates for its continuous-in-time formulation using general order polynomials are found in e.g. Elliott and Ranner [2021]. For the sake of generality, here we focus on moving surfaces with boundary. The problem reads: Find $u_{h}\in V_{h}(\Gamma_{h})$ such that
$\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+\overline{a}_{h}(u_{h}\,,\,v_{h})+(({\boldsymbol{b}}_{rel}u_{h}-\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}u_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}}\,,\,v_{h})_{\partial\Gamma_{h}}$ (38)
@@
$\partial\Gamma_{h,D}\cap\partial\Gamma_{h,N}=\emptyset,\quad\partial\Gamma_{h,D}\cup\partial\Gamma_{h,N}=\partial\Gamma_{h}.$ (40)
Nitsche’s boundary penalty method is used in order to impose the inhomogeneous Dirichlet boundary condition $u_{h}(t,{\boldsymbol{x}})|_{\partial\Gamma_{h,D}}=u(t,{\boldsymbol{x}})$ weakly for the diffusive term Nitsche [1971], Ern and Guermond [2021a]. For convenience, we define $v^{+}=\max\{v,0\}$ and $v^{-}=\min\{v,0\}$. The resulting bilinear forms are given by
$a_{h}(u_{h}\,,\,v_{h})=$ $~\overline{a}_{h}(u_{h}\,,\,v_{h})-((\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}u_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}}\,,\,v_{h})_{\partial\Gamma_{h,D}}$ (41)
@@
$\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+a_{h}(u_{h}\,,\,v_{h})=l_{h}(v_{h}),\;\forall v_{h}\in V_{h}(\Gamma_{h}).$ (42)
@@ 5.1 Numerical methods for advection-dominant problems
In biophysical cell models, as introduced in subsection 1.1, it is not uncommon to incur advection-dominant ADR equations. In our case, even purely parabolic equations might become advection-dominant due to the introduction of an ALE velocity ${\boldsymbol{v}}^{\mathcal{A}}$. In the present work we employ and extend the continuous interior penalty (CIP) stabilization to cope with the possibly dominant advection regime. The CIP method was proposed and analyzed in Burman and Hansbo [2004], Burman and Fernández [2009] and then later extended to the case of stationary surfaces in Burman et al. [2019]. Moreover, the CutFEM approach for moving surfaces developed in Hansbo et al. [2015] applies a CIP-type stabilization for advection dominant problems on moving surfaces. We chose CIP for the following reasons:
@@
  • Although only weakly consistent, it has proven to be very successful in time-dependent problems given its commutativity with the time derivative Ern and Guermond [2021c];
@@
$\{f\}|_{F}=\frac{1}{2}\left(f_{F}^{+}+f_{F}^{-}\right),\quad{\llbracket f\rrbracket|_{F}}=f_{F}^{+}-f_{F}^{-},$ (43)
@@
$s_{h}^{b}\left({u}_{h},{v}_{h}\right):=\gamma_{b}\sum_{F\in\mathcal{F}_{h}^{i}}\phi_{F}h_{F}\left(\llbracket{\boldsymbol{b}_{M_{h}}}\cdot\nabla_{M_{h}}{u}_{h}\rrbracket,\llbracket\boldsymbol{b}_{M_{h}}\cdot\nabla_{M_{h}}{v}_{h}\rrbracket\right)_{F},$ (44)
@@
$\phi_{F}=\max(\phi_{T^{+}},\phi_{T^{-}}),\quad\phi_{T^{\pm}}=\min(\beta_{T^{\pm}}^{-1}h_{T^{\pm}},\mu_{T^{\pm}}^{-1}),\quad T^{\pm}\in\mathcal{T}_{h}.$ (45)
In the above $\beta_{T}$ is a local velocity scale and $\mu_{T}$ is the reciprocal of a time Ern and Guermond [2021c]. If ${\boldsymbol{b}}_{M_{h}}\in L^{\infty}(M_{h})\cap C^{0}(M_{h})$, then (44) can be simplified to
$s_{h}^{b}({u}_{h},{v}_{h}):=\gamma_{b}\sum_{F\in\mathcal{F}_{h}^{i}}\beta_{F}h_{F}^{2}\left(\llbracket\nabla_{M_{h}}{u}_{h}\rrbracket,\llbracket\nabla_{M_{h}}{v}_{h}\rrbracket\right)_{F}.$ (46)
@@
$\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+a_{h}(u_{h}\,,\,v_{h})+s_{h}^{b}(u_{h}\,,\,v_{h})=l_{h}(v_{h}),\;\forall v_{h}\in V_{h}(\Gamma_{h}).$ (47)
@@ 5.2 Numerical results for the ADR system
$\mathbb{A}(t)=\begin{bmatrix}\cos(t)&-\sin(t)&0\\ \sin(t)&\cos(t)&0\\ 0&0&1\end{bmatrix},\quad B(t)=\begin{bmatrix}0\\ 0\\ 0\end{bmatrix}$. (48)
@@
$\partial_{t}u+\nabla_{\Gamma}\cdot({\boldsymbol{b}}_{\Gamma}u)+c_{\Gamma}u=f,$ (49)
@@
${\boldsymbol{b}}_{\Gamma}$ $={\boldsymbol{v}}^{\mathcal{A}}+{\boldsymbol{b}}_{rel}={\boldsymbol{v}}^{\mathcal{A}}+\mathbb{P}_{\Gamma}(2,-x_{1},0),$
@@
    adrSolver 1: CIP stabilized solver (47).
@@
To further test the effectiveness of the algorithm we now consider an ill-posed problem. Maintaining the half-sphere geometry and the transformation as in (48), we solve the pure transport problem
$\partial_{t}u+\nabla_{\Gamma}({\boldsymbol{b}}_{\Gamma}u)=0,$ (51)
@@
${\boldsymbol{b}}_{\Gamma}={\boldsymbol{v}}^{\mathcal{A}}+{\boldsymbol{b}}_{rel}={\boldsymbol{v}}^{\mathcal{A}}+(x_{3},0,-x_{1})\cdot(1-e^{-10x_{3}}),\quad u_{0}=e^{-3(x_{1}^{2}+x_{2}^{2})}.$ (52)
@@
It can be seen that the non-stabilized solver is unstable due to the pure advective regime, leading to oscillations that inevitably effect the whole domain, see Figure 4(a). adrSolver 1 effectively controls the solution’s gradient thanks to the stabilization, generating a non-oscillating solution, but fails to maintain its positivity, see Figure 5 (right plot). This characteristic is lost along $\partial\Gamma$ where the species encounters the steep boundary layer. adrSolver 2 manages, thanks to the cutoff, to keep the solution positive but fails in maintaining mass-preservation as it can be seen in Figure 5 (left plot), where the relative absolute percentage mass error is plotted. adrSolver 4 successfully manages to deal with the boundary layer, maintain positivity, and avoid oscillations. The final solution is shown in Figure 4(b).
@@ 6 Cahn-Hilliard system as a phase separation model
Phase field modeling plays a central role in biophysical systems to investigate phase transition processes. Common phase field models for stationary surfaces are derived from Allen-Cahn and Cahn-Hilliard dynamics Cahn and Hilliard [1958], Erlebacher et al. [2001], Elliott and Stinner [2010], Du et al. [2011], where the motion follows the minimization of an energy functional. We will focus here on the functional
$\mathcal{E}_{CH}(u)=\int_{M(t)}\sigma\Bigl(\underbrace{\frac{\varepsilon|\nabla_{\Gamma}u|^{2}}{2}}_{I}+\underbrace{\frac{1}{\varepsilon}F(u)}_{II}\Bigr),$ (53)
@@
$F_{1}(u)$ $:=\frac{1}{4}((1-u)\log(1-u)+(1+u)\log(1+u))+\frac{1-u^{2}}{2},$ (54a)
$F_{2}(u)$ $:=\frac{(1-u^{2})^{2}}{4},$ (54b)
@@
$F_{1}^{\prime}(u)=\frac{1}{4}\log\Bigl(\frac{1+u}{1-u}\Bigr)-u,\text{ and }F_{2}^{\prime}(u)=u^{3}-u\text{ respectively}.$ (55)
$F_{1}$ is a Flory-Huggins-type mixing free energy density whose formulation finds its roots in configurational entropy Huggins [1941], Flory [1942]. It can be seen that $F_{1}$ is only well posed for $u\in(-1,1)$. For this reason, a simpler polynomial potential $F_{2}$ is often used. The advantage of the latter is that it is well defined for values of $u$ outside the interval (1,1)(-1,1).
The extension of the Cahn-Hilliard equation to evolving surfaces has seen recent developments due to its applicability to biological membranes and alloys Zhiliakov et al. [2021], Eilks and Elliott [2008]. The evolving surface Cahn-Hillard equation Caetano and Elliott [2021], Caetano et al. [2023] can be written as
$\begin{cases}\partial^{\bullet}u+(\nabla_{\Gamma}\cdot{\boldsymbol{v}})u=\nabla_{\Gamma}(m\nabla_{\Gamma}w),\\ w=-\sigma\varepsilon\Delta_{\Gamma}u+\frac{\sigma}{\varepsilon}F^{\prime}(u),\end{cases}$ (56)
subject to the initial condition $u(0)=u_{0}$ for suitable initial data. $w$ represents the chemical potential and $m\in\mathbb{R}^{+}$ is the mobility. As explained in Liu et al. [2024], models for evolving surfaces have similar energy functions to models on stationary surfaces, but important differences arise. Even if the evolution of the systems is driven by the minimization of (53), energy dissipation is no longer satisfied since the velocity is viewed as external force that adds to the system. Analogously, mass preservation depends on the velocity. To keep mass conservation, we must assume $\nabla_{\Gamma}\cdot{\boldsymbol{v}}=0$, which is equivalent to assume inextensibility of the cell membrane, and by consequence area preservation. This assumption also allows to consistently extend this one-phase model to $N$-phase models satisfying the hyperlink condition, as described in Liu et al. [2024].
@@ 6.1 Numerical methods for Cahn-Hilliard system
For the discretization of (56), the evolutionary surface finite element method (ESFEM) is one of the most popular methods adopted and the one we consider here Dziuk and Elliott [2013]. Its advantage lies in simple implementation, low memory consumption, ability to handle complex deformations and potential for efficient parallelization. We refer to Liu et al. [2024] for a review of alternative methods. We employ a semi-implicit approach where the fully discrete formulation reads: given $u_{h}^{n}\in V_{h}(\Gamma_{h}^{n})$ find $u^{n+1}_{h},w^{n+1}_{h}\in V_{h}(\Gamma_{h}^{n})$ such that
$n_{h}\mathopen{}\left(\frac{u^{n+1}_{h}-u_{h}^{n}}{\tau}\,,\,\phi_{h}\mathclose{}\right)+(({\boldsymbol{v}}_{h}-{\boldsymbol{v}}^{\mathcal{A}}_{h})u^{n+1}_{h}\,,\,\nabla_{\Gamma^{n}_{h}}\phi_{h})_{\Gamma^{n}_{h}}+mb_{h}(w^{n+1}_{h}\,,\,\phi_{h})=0,$ (57a)
$n_{h}(w^{n+1}_{h}\,,\,\psi_{h})-\sigma\varepsilon b_{h}(u^{n+1}_{h}\,,\,\psi_{h})=\frac{\sigma}{\varepsilon}n_{h}(F^{\prime}(u^{n}_{h})\,,\,\psi_{h}),$ (57b)
@@
$n_{h}(u_{h}\,,\,\phi_{h})=(u_{h}\,,\,\phi_{h})_{\Gamma_{h}^{n}},\quad b_{h}(u_{h}\,,\,\psi_{h})=(\nabla_{\Gamma_{h}^{n}}u_{h}\,,\,\nabla_{\Gamma_{h}^{n}}\psi_{h})_{\Gamma_{h}^{n}}.$ (58)
Similar discretizations can be found in Elliott and Stinner [2010], Bachini et al. [2023], Mokbel et al. [2024], where a Cahn-Hilliard solver is coupled with a domain evolution which is unknown. In those works, the potential $F_{2}$ was used. In our case, we can take advantage of the scheme presented in Section 3 to employ both $F_{1}$ and $F_{2}$, and apply mass conservation if needed. The scheme proceeds as follows.
In the predictor step we solve the unconstrained problem (57). Taking the non-linear term explicitly ensures $u_{h}^{n}$ lays inside the desired bound.